# Evaluation-Aware Reinforcement Learning

### v1

@@ 1 Introduction
Policy evaluation is often a prerequisite for deploying RL policies in safety- or performance-critical settings such as industrial control (Gao et al., 2014; Zheng et al., 2021), robotics (Andrychowicz et al., 2020; Kalashnikov et al., 2018) and healthcare (Komorowski et al., 2018; Fox et al., 2021). Evaluation settings are often not representative of deployment settings, due to factors such as data availability, data collection costs, and safety constraints. For example, consider a self-driving car that cannot be exhaustively tested in every possible real-world scenario that it will encounter; instead, it may be tested via a combination of limited tests in the real world (perhaps with professional drivers ready to take over), controlled scenarios such as a test course, tests in simulation, or off-policy evaluation via historical data. More generally, compared to the setting that a policy will be deployed in, evaluation data may be collected in a manner that is off-policy, limited in size, has limited support, or uses an inaccurate model of the environment.
@@ 2 Background and Related Work
Policy Evaluation. Policy evaluation refers to estimating $V_{\pi}(s)$ or its start-state expectation $J_{\pi}$. The simplest approach is Monte Carlo (MC) evaluation, where trajectories are sampled on-policy and averaged as $\hat{J}_{\pi}^{MC}\coloneqq\tfrac{1}{N}\sum_{i=1}^{N}g(h_{i})$. Since on-policy evaluation can be costly, off-policy policy evaluation (OPE) methods estimate policy performance from historical data (Uehara et al., 2022). The standard importance sampling (IS) estimator reweights trajectories from a behavior policy $\pi_{b}$, given by $\hat{J}_{\pi}^{IS}\coloneqq\tfrac{1}{N}\sum_{i=1}^{N}g(h_{i})\prod_{t=0}^{T^{i}-1}\tfrac{\pi(a^{i}_{t}|s^{i}_{t})}{\pi_{b}(a^{i}_{t}|s^{i}_{t})}$. Per-decision importance sampling (PDIS) (Precup et al., 2000) applies the weights step by step, leading to $\hat{J}_{\pi}^{PDIS}\coloneqq\tfrac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{T^{i}-1}\gamma^{t}r_{t}^{i}\prod_{k=0}^{t}\tfrac{\pi(a_{k}^{i}|s_{k}^{i})}{\pi_{b}(a_{k}^{i}|s_{k}^{i})}$, which reduces variance compared to trajectory-level IS. Fitted Q-Evaluation (FQE) (Le et al., 2019) instead learns the Q-function of the evaluation policy via Bellman consistency and derives $J_{\pi}$ from the learned value function. Doubly robust (DR) estimators (Jiang & Li, 2016; Thomas & Brunskill, 2016) combine PDIS with approximate value functions $\hat{Q}$ and $\hat{V}$; a common formulation is $\hat{J}_{\pi}^{DR}\coloneqq\tfrac{1}{N}\sum_{i=1}^{N}\big[\hat{V}(s_{0}^{i})+\sum_{t=0}^{T^{i}-1}\gamma^{t}w_{t}^{i}\big(r_{t}^{i}+\gamma\hat{V}(s_{t+1}^{i})-\hat{Q}(s_{t}^{i},a_{t}^{i})\big)\big]$, where $w_{t}^{i}=\prod_{k=0}^{t}\tfrac{\pi(a_{k}^{i}|s_{k}^{i})}{\pi_{b}(a_{k}^{i}|s_{k}^{i})}$. DR estimators reduce variance while maintaining consistency if either the importance weights or the value approximations are accurate. For further discussion, we refer to the survey by Uehara et al. (2022). In this work, we compare against FQE, TIS, PDIS, and DR.
@@ 3 Evaluation-Aware Reinforcement Learning
With $\hat{V}_{D}^{\pi}$ estimated from assessment rollouts, we use deployment interactions to approximate the target state-value function $V_{D}^{\pi}$ for policy learning. In a policy-gradient setting, for example, Monte Carlo return samples starting from state $s$ serve as target values $V_{D}^{\pi}(s)$, as in standard on-policy policy gradient algorithms such as A2C (Konda & Tsitsiklis, 1999) and TRPO (Schulman et al., 2015).
@@ 3.1 Theoretical insights into EvA-RL
For a detailed proof, see Appendix A.1. Briefly, the optimization problem can be shown to be a quadratically constrained linear program (Boyd & Vandenberghe, 2004) with a positive semi-definite quadratic constraint. ∎
Thus, multiple optimal solutions for $V_{D}^{\pi}$ may exist for the Bellman-relaxed problem 5, each corresponding to a distinct policy. These solutions differ from each other by vectors in the null-space of the quadratic constraint. Incorporating Bellman-consistency constraints in the optimization may still admit multiple optimal $V_{D}^{\pi}$. This is in contrast to standard RL where the optimal value function is unique (Sutton & Barto, 1998). Additionally, the expected return under Bellman-relaxed optimization forms an upper bound on the expected return under Bellman consistency. We provide the exact expression for the upper bound in Appendix A.3.
@@ 5 Conclusion
6 Acknowledgement
This work has taken place in the Safe, Correct, and Aligned Learning and Robotics Lab (SCALAR) at The University of Massachusetts Amherst. SCALAR research is supported in part by the NSF (IIS-2437426), the Long-Term Future Fund, and Open Philanthropy.
@@ A.2 Analysis of the optimization problem
Therefore, the problem is an instance of convex quadratic program (QP) over state-values (Boyd & Vandenberghe, 2004). It can further be categorized as a quadratically constrained linear program (QCLP). ∎

### v2

@@ 1 Introduction
Policy evaluation is often a prerequisite for deploying RL policies in safety- or performance-critical settings such as industrial control (Mirhoseini et al., 2020), robotics (Andrychowicz et al., 2020; Kalashnikov et al., 2018) and healthcare (Fox et al., 2020). Evaluation settings are often not representative of deployment settings, due to factors such as data availability, data collection costs, and safety constraints. For example, consider a self-driving car that cannot be exhaustively tested in every possible real-world scenario that it will encounter; instead, it may be tested via a combination of limited tests in the real world (perhaps with professional drivers ready to take over), controlled scenarios such as a test course, tests in simulation, or off-policy evaluation via historical data. More generally, compared to the setting that a policy will be deployed in, evaluation data may be collected in a manner that is off-policy, limited in size, has limited support, or uses an inaccurate model of the environment.
@@ 2 Background and Related Work
Policy Evaluation. Policy evaluation refers to estimating $V_{\pi}(s)$ or its start-state expectation $J_{\pi}$. The simplest approach is Monte Carlo (MC) evaluation, where trajectories are sampled on-policy and averaged as $\hat{J}_{\pi}^{MC}\coloneqq\tfrac{1}{N}\sum_{i=1}^{N}g(h_{i})$. Since on-policy evaluation can be costly, off-policy policy evaluation (OPE) methods estimate policy performance from historical data (Uehara et al., 2022). The standard importance sampling (IS) estimator reweights trajectories from a behavior policy $\pi_{b}$, given by $\hat{J}_{\pi}^{IS}\coloneqq\tfrac{1}{N}\sum_{i=1}^{N}g(h_{i})\prod_{t=0}^{T^{i}-1}\tfrac{\pi(a^{i}_{t}|s^{i}_{t})}{\pi_{b}(a^{i}_{t}|s^{i}_{t})}$. Per-decision importance sampling (PDIS) (Precup et al., 2000) applies the weights step by step, leading to $\hat{J}_{\pi}^{PDIS}\coloneqq\tfrac{1}{N}\sum_{i=1}^{N}\sum_{t=0}^{T^{i}-1}\gamma^{t}r_{t}^{i}\prod_{k=0}^{t}\tfrac{\pi(a_{k}^{i}|s_{k}^{i})}{\pi_{b}(a_{k}^{i}|s_{k}^{i})}$, which reduces variance compared to trajectory-level IS. Fitted Q-Evaluation (FQE) (Le et al., 2019) instead learns the Q-function of the evaluation policy via Bellman consistency and derives $J_{\pi}$ from the learned value function. Doubly robust (DR) estimators (Jiang and Li, 2016; Thomas and Brunskill, 2016) combine PDIS with approximate value functions $\hat{Q}$ and $\hat{V}$; a common formulation is $\hat{J}_{\pi}^{DR}\coloneqq\tfrac{1}{N}\sum_{i=1}^{N}\big[\hat{V}(s_{0}^{i})+\sum_{t=0}^{T^{i}-1}\gamma^{t}w_{t}^{i}\big(r_{t}^{i}+\gamma\hat{V}(s_{t+1}^{i})-\hat{Q}(s_{t}^{i},a_{t}^{i})\big)\big]$, where $w_{t}^{i}=\prod_{k=0}^{t}\tfrac{\pi(a_{k}^{i}|s_{k}^{i})}{\pi_{b}(a_{k}^{i}|s_{k}^{i})}$. DR estimators reduce variance while maintaining consistency if either the importance weights or the value approximations are accurate. For further discussion, we refer to the survey by Uehara et al. (2022). In this work, we compare against FQE, TIS, PDIS, and DR.
@@ 3 Evaluation-Aware Reinforcement Learning
With $\hat{V}_{D}^{\pi}$ estimated from assessment rollouts, we use deployment interactions to approximate the target state-value function $V_{D}^{\pi}$ for policy learning. In a policy-gradient setting, for example, Monte Carlo return samples starting from state $s$ serve as target values $V_{D}^{\pi}(s)$, as in standard on-policy policy gradient algorithms such as A2C (Konda and Tsitsiklis, 1999) and TRPO (Schulman et al., 2015).
@@ 3.1 Theoretical insights into EvA-RL
For a detailed proof, see Appendix A.1. Briefly, the optimization problem can be shown to be a quadratically constrained linear program (Boyd and Vandenberghe, 2004) with a positive semi-definite quadratic constraint. ∎
Thus, multiple optimal solutions for $V_{D}^{\pi}$ may exist for the Bellman-relaxed problem 5, each corresponding to a distinct policy. These solutions differ from each other by vectors in the null-space of the quadratic constraint. Incorporating Bellman-consistency constraints in the optimization may still admit multiple optimal $V_{D}^{\pi}$. This is in contrast to standard RL where the optimal value function is unique (Sutton and Barto, 1998). Additionally, the expected return under Bellman-relaxed optimization forms an upper bound on the expected return under Bellman consistency. We provide the exact expression for the upper bound in Appendix A.3.
@@ A.2 Analysis of the optimization problem
Therefore, the problem is an instance of convex quadratic program (QP) over state-values (Boyd and Vandenberghe, 2004). It can further be categorized as a quadratically constrained linear program (QCLP). ∎